OpenStudy (anonymous):

Please show steps so I understand this thoroughly.. (: Thanks! Find the product. (x - nu)(9x + 3nu)

OpenStudy (ipwnbunnies):

Do you know the FOIL method?

OpenStudy (anonymous):

hey ipwnbunnies

OpenStudy (anonymous):

Not fairly well, however I could probably pickup on it.. If you guide me through it.

OpenStudy (ipwnbunnies):

Have you done any quadratic equation stuff? Like, if you had (x+2)(x+3), can you expand it out?

OpenStudy (anonymous):

Wouldn't that be x^2+5?

OpenStudy (ipwnbunnies):

Not quite. It'll be x^2 + 5x + 6. Hold on.

OpenStudy (ipwnbunnies):

Check out this little lesson: http://www.mathsisfun.com/definitions/foil-method.html

OpenStudy (anonymous):

follow the distributive law

OpenStudy (anonymous):

The page is blank for me @iPwnBunnies =[

OpenStudy (ipwnbunnies):

Try this one lol: http://www.freemathhelp.com/using-foil.html

OpenStudy (anonymous):

9x^2+3xnu-9xnu-3n^2x^2 9x^2-6xnu-3n6@u^2 i doubled check and this is correct

OpenStudy (anonymous):

9x^2-6xnu-3n^2u^2

OpenStudy (anonymous):

i tripled checked to make sure

OpenStudy (anonymous):

Okay, let me check to see if I do this right though... So.. (x - nu)(9x + 3nu) x * 9x = 9x^2 x * 3nu = 3nux ? -nu * 9x = -9nux -nu * 3nu = -3nu Is that all right?

OpenStudy (ipwnbunnies):

No, I'm wrong. Check your last product.

OpenStudy (anonymous):

no -nu*3nu=-3(nu)^2

OpenStudy (anonymous):

Oh .. Whoops lol.

OpenStudy (anonymous):

So now.. 9x^2 + 3nux - 9 nux + -3nu^2 9x^2 -6nux - 3nu^2 ?? Right

OpenStudy (anonymous):

no -3n^2u^2 thats why i used parenthesis last time. to tell you that both the n and the u are squared

OpenStudy (anonymous):

Oh. Ok. Are they correct @iPwnBunnies ?

OpenStudy (ipwnbunnies):

What Bsho posted is correct. You just messed up the 'L' in FOIL. the last term should be -3*n^2*u^2

OpenStudy (ipwnbunnies):

Now, add up the terms to get your answer.

OpenStudy (anonymous):

So instead it would be 9x^2 - 6nux -3n^2u^2 ?

OpenStudy (anonymous):

that is correct

OpenStudy (ipwnbunnies):

Correct, good job.

OpenStudy (anonymous):

Thank you @iPwnBunnies and @Bsho1997 :)

OpenStudy (anonymous):

your welcome